OpenStudy (anonymous):

Find the derivative: 2ysin^2(xy) Thanks, Find the derivative: 2ysin^2(xy) Thanks, @Mathematics

OpenStudy (anonymous):

i assume this is set equal to something, or are you looking for a partial?

OpenStudy (anonymous):

Its part of a bigger function, but I'm just confused on how to find the derivative of this,

OpenStudy (anonymous):

you can't unless it is equal something.

OpenStudy (anonymous):

oh my bad, so the function is basically 2ysin^2(xy) + 6x = 3 + (pi)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\[2(y'\sin^2(xy)+2y\cos(xy)y'+6=0\]

OpenStudy (anonymous):

solve for y', and the 2 is in front of the first part, not the 6

OpenStudy (anonymous):

So would i use the chain rule for ysin^2 and not the (xy)?

OpenStudy (anonymous):

\[2(y'\sin^2(xy)+2y\cos(xy)y')+6=0\] without an equation you cannot solve. yes i used the chain rule for both

OpenStudy (anonymous):

oh damn good catch

OpenStudy (anonymous):

\[2(y'\sin^2(xy)+2y\sin(xy)\cos(xy)y')+6=0\] is more like it. my mistake, good call

OpenStudy (anonymous):

Oh i see, thanks a lot, I appreciate it.

OpenStudy (anonymous):

yw, and hope i didn't confuse with the wrong solution!

OpenStudy (anonymous):

not at all, haha thanks.

OpenStudy (anonymous):

haha?

OpenStudy (anonymous):

Whats wrong with typing haha? I was just trying to be friendly.